OpenStudy (anonymous):

Find the length of the curve y= x⁴/4 + 1/4x² from x= 1 to x = 2

sam (.sam.):

\[y=\frac{x^4}{4}+\frac{x^2}{4}\]

sam (.sam.):

Do you know how to get this? \[\int\limits_1^2 \sqrt{\left(x^3+\frac{x}{2}\right)^2+1} \, dx\]

OpenStudy (anonymous):

yes...thank you